A Toolkit for Algebraic Geometry

Ashutosh Rege ^{∗}

Computer Science Division University of California Berkeley, CA 94720

Abstract

In this talk we describe a toolkit for polynomial algebra developed at Berkeley for solving various problems in algebraic geometry and polynomial system solving. The toolkit can be applied to a number of problems arising in robotics, computer-aided geometric design, and other areas of science and engineering. The toolkit uses various techniques like straight-line programs and mixed arithmetic to make multi-variate polynomial computations more efficient. We demonstrate the application of this toolkit to the well known problem of geometric theorem proving and to solving other problems.

Toolkit availability : Version 1.1 of the toolkit can be obtained by anonymous ftp from robotics.eecs.berkeley.edu.

∗E-mail: rege@cs.Berkeley.edu